OpenStudy (hdrager):

f(x)=sin(2x) f(pi/4)=?

OpenStudy (hdrager):

I still don't know what to do???

OpenStudy (legomyego180):

You need to have the unit circle memorized to do this problem if you havent already memorized it. \[f(x)=\sin(2x)\] \[f(\frac{ \pi }{ 4 })=\sin(2*\frac{ \pi }{ 4 })=\sin(\frac{ 2\pi }{ 4 })=\sin(\frac{ \pi }{ 2 })\]

OpenStudy (legomyego180):

Then take a look at your unit circle... https://www.google.com/search?q=unit+circle&source=lnms&tbm=isch&sa=X&ved=0ahUKEwiz7cjp_cHOAhUN02MKHatPCjsQ_AUICCgB&biw=1440&bih=791#imgrc=7btvgKywl0C4yM%3A Look at the angle \(\frac{ \pi }{ 2 }\) and the sin value and you'll see that \(\sin(\frac{ \pi }{ 2 })=1\)

OpenStudy (hdrager):

ohhh alright

OpenStudy (legomyego180):

memorize it asap! I waited forever to memorize it and had to end up retaking trig. It's extremely important

OpenStudy (hdrager):

I almost had it memorized last year but I've forgotten over the summer break